A potential field is given by V 5 3x2 y 2 yz. Which of the following is not true?


1
Expert's answer
2021-06-30T10:13:53-0400

Gives

"V=3x^2y^2-yz"

"E=-\nabla V"

"E=-(\frac{dV}{dx}\hat{a_x}+\frac{dV}{dy}\hat{a_y}+\frac{dV}{dz}\hat{a_z})"

"E_x=-\frac{dV}{dx}=6xy^2"

"E_y=\frac{dV}{dy}=6x^2y-z"

"E_z=\frac{dV}{dz}=-y"


"E=-(6xy^2\hat{a_x}+(6x^2y-z)\hat{a_y}-y\hat{a_z})"